OpenStudy (anonymous):

Factor the expression x2 − 5x − 24.

OpenStudy (anonymous):

x^2 + 3x - 8x - 24 = (x + 3)(x - 8) First you should find two numbers whose sum is -5 and product is -24(in this case 3 and -8). Then you should split the middle term like I have done here. Group the terms and factor - there you have the factored expression!!

OpenStudy (amorfide):

do the coefficient of x² multiplied by the constant 1x-24=-24 what two numbers multiply to give -24 but add to give -5? list factors of -24 1 24 2 12 3 8 4 6 -1 24 -2 12 -3 8 -4 6 1 -24 2 -12 3 -8 4 -6 3+-8=-5 use that you now have (x + 3 ) (x - 8)